OpenStudy (anonymous):

How do we transform System A into System B? System A 3x-5y=3 7x-8y=-15 System B 3x-5y=3 x+2y=-21 I just can't figure out what I multiple 7x-8y=-15 by to get x+2y=-21.

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (amoodarya):

multiply first equation by -2 then add by 2nd equation you will have the answer easy

OpenStudy (amoodarya):

3x-5y=3 7x-8y=-15 multiply by -2 the 1st eq. -6x+10y=-6 so add by 7x-8y=-15

OpenStudy (amoodarya):

3x-5y=3 7x-8y=-15 3x-5y=3 (-6x+10y=-6 ) +(7x-8y=-15 )

OpenStudy (anonymous):

Oh, I was trying to do it completely wrong. Thank you!